When the handle on a car jack is pushed through 40 cm, the 10,000-N car is lifted 0.8 cm' Find (a) the work done on the car, (b) the force exerted on thejack handle, assuming no losses, and (c) the mechanical

advantageMA = Fs1,7/Fn, and show that this has the expected value'
Answer #1

(a) Work = force * distance = 10000* 0.8/100 = 80 J

(b) work = force * distance

80 = force * 40/100

force = 200 N

(c) MA = 10000/200 = 50